# Malicious Code Detection in Smart Contracts via Opcode Vectorization

### v1

@@ 1. Introduction
With the rapid development of blockchain technology, smart contracts have seen increasingly widespread applications in finance, supply chains, insurance, and other fields. A smart contract is a computer protocol that automatically executes predefined conditions and logic. It is essentially a piece of code designed to facilitate, verify, or enforce the terms of a contract (Sayeed et al., 2020). Although smart contracts offer advantages such as decentralization, transparency, and immutability, they still face security risks. In 2017, the Parity Multisig wallet had a fatal unauthorized access vulnerability, resulting in \$300 million being frozen (Metz, 2016). In 2018, an integer overflow vulnerability in the BEC campaign caused the instantaneous evaporation of over \$900 million. Security incidents in blockchain can be categorized into the following types: trading platform security incidents, user account and key security incidents, miner node system security incidents, smart contract security incidents, and other security incidents. Among these, smart contract security incidents account for approximately 6%, but the losses incurred are substantial, accumulating to \$5.1 billion by the end of 2022. Therefore, addressing smart contract security issues is urgent (Chen et al., 2020).
Traditional static and dynamic analysis methods face certain limitations in detecting malicious code in smart contracts, including low analysis efficiency, high false-positive rates, and an inability to identify new and complex vulnerabilities (Tchakounté et al., 2022). In recent years, machine learning technology has achieved remarkable results in fields such as computer vision and natural language processing, opening up new research directions for detecting malicious code in smart contracts.
Machine learning-based malicious code detection technology for smart contracts can more effectively identify and prevent vulnerabilities and malicious behaviors in smart contracts (Dwivedi et al., 2021). It enables automated and intelligent auditing of smart contract code, reducing the cost of manual audits while improving audit efficiency and accuracy (Ajienka et al., 2020). In the future, as research deepens and technology advances, machine learning-based malicious code detection technology is certain to play an even greater role in the field of smart contracts.
@@ 2. Background
Smart contracts are self-executing and self-verifying computer protocols designed to facilitate, verify, or enforce the negotiation of contracts between parties (Barboni et al., 2023). Smart contracts enable trustworthy transactions without third-party intermediaries, thereby reducing the risks of fraud and default. The concept of smart contracts was first proposed by computer scientist Nick Szabo in 1994 (Ghaleb et al., 2023).
Smart contracts are one of the core applications of blockchain technology, implemented by deploying programmable scripts on the blockchain (Sharma et al., 2023). These scripts can automatically trigger, execute, and verify relevant operations when specific conditions are met (Tolmach et al., 2021). Since smart contracts run on decentralized blockchain networks, they have the following characteristics:
@@ 3.1. Feature Extraction
The features obtained from Ethereum smart contracts cannot be directly used as machine learning inputs, thus vectorization of smart contract features becomes a necessary step (Wood et al., 2014). This paper employs an opcode-based text vectorization method that first categorizes opcodes according to their semantic meanings in the Ethereum Virtual Machine (EVM), then transforms the simplified opcodes into vectors using N-Gram and TF-IDF algorithms.
@@ 3.2. Problem Analysis
In recent years, machine learning-based smart contract vulnerability detection techniques have made significant progress. These methods analyze smart contract source code, bytecode, and opcodes to better detect potential vulnerabilities (Zhuang et al., 2020).
@@ 4.1. Dataset Collection
Etherscan (Tan Matai, [n. d.]) is an Ethereum blockchain explorer and analytics platform. It allows users to query and retrieve information about transactions, addresses, blocks, tokens (including ERC-20 and ERC-721 tokens), and smart contracts on the Ethereum network. Etherscan provides developers, researchers, and other participants in the Ethereum ecosystem with a user-friendly interface to track and analyze activities on the Ethereum blockchain.
@@ 4.2. Label Annotation
We used the Ethereum static analysis tool Slither (Feist et al., 2019) to detect vulnerabilities in smart contracts. After configuring the Slither tool and selecting the compiler version (solc) according to the pragma solidity ^0.6.0 statement in the smart contract code, we performed the detection. The detection results for one sample smart contract (as shown in Figure 4) revealed both access control vulnerabilities and reentrancy attack vulnerabilities.
@@
[…]
@@ 6. Related Work
In machine learning-based malicious code detection, feature extraction and representation are crucial (Liu and Li, 2025). Researchers attempt to extract various features from smart contract code, including syntactic features, semantic features, control flow, and data flow (Bu et al., 2025b). These features can be represented using techniques such as the bag-of-words model, word vector embeddings, and graph representation learning (Asudani et al., 2023). Recent studies further demonstrate the importance of transaction pattern analysis, as seen in (Li et al., 2024e), which detects malicious Web3 accounts through transaction graph analysis. Effective feature extraction and representation methods help improve the detection performance of machine learning models (Li et al., 2024b).
@@
In 2021, Xu et al. (Xu et al., 2021) extracted corresponding abstract syntax trees from the source code of each vulnerability dataset, as well as from a large dataset. They then extracted feature vectors based on subnodes shared by the two ASTs. Finally, they used the K-Nearest Neighbor (KNN) algorithm and Stochastic Gradient Descent (SGD) to train models. Both methods leveraged abstract syntax trees as the basis for feature extraction, avoiding the need to set up complex patterns or execute them on Ethereum. Recent advancements like (Li et al., 2024c) and (Li et al., 2024d) further extend this paradigm by detecting state defects in decentralized exchanges and analyzing cross-contract execution patterns. This makes these methods relatively simple for professionals to use. However, a limitation of these approaches is that obtaining source code is more difficult compared to other contract attributes, reducing their practical utility for third-party detection. Alternative solutions like (Li et al., 2020) and (Liu et al., 2024b) address this challenge by characterizing account behaviors and detecting malicious activities through gas usage patterns.
@@
Research on machine learning-based malicious code detection in smart contracts has made progress but still faces many challenges and problems. The integration of AI and blockchain for privacy preservation (Li et al., 2023), the detection of emerging threats like NFT wash trading (Niu et al., 2024), and the need to characterize complex ecosystems like Solana NFTs (Kong et al., 2024) highlight the evolving nature of this field. Additionally, emerging tools like (Wang et al., 2024) statistically quantify external data dependencies in real-world contracts, while (Li et al., 2025) demonstrates how LLMs can automate bad practice detection, pointing to future research directions.

### v2

@@ 1. Introduction
With the rapid development of blockchain technology, smart contracts have seen increasingly widespread applications in finance, supply chains, insurance, and other fields. A smart contract is a computer protocol that automatically executes predefined conditions and logic. It is essentially a piece of code designed to facilitate, verify, or enforce the terms of a contract (Sayeed et al., 2020). Although smart contracts offer advantages such as decentralization, transparency, and immutability, they still face security risks. In 2017, the Parity Multisig wallet had a fatal unauthorized access vulnerability, resulting in \$300 million being frozen (Metz, 2016). In 2018, an integer overflow vulnerability in the BEC campaign caused the instantaneous evaporation of over \$900 million. Security incidents in blockchain can be categorized into the following types: trading platform security incidents, user account and key security incidents, miner node system security incidents, smart contract security incidents, and other security incidents. Among these, smart contract security incidents account for approximately 6%, but the losses incurred are substantial, accumulating to \$5.1 billion by the end of 2022. Therefore, addressing smart contract security issues is urgent (Chen et al., 2020; Jiao et al., 2024; Kumar et al., 2024; Wei et al., 2024; Zhu et al., 2024).
Traditional static and dynamic analysis methods face certain limitations in detecting malicious code in smart contracts, including low analysis efficiency, high false-positive rates, and an inability to identify new and complex vulnerabilities (Tchakounté et al., 2022; Wang and Zhao, 2024; Grossman et al., 2024; Wang et al., 2024a). In recent years, machine learning technology has achieved remarkable results in fields such as computer vision and natural language processing, opening up new research directions for detecting malicious code in smart contracts.
Machine learning-based malicious code detection technology for smart contracts can more effectively identify and prevent vulnerabilities and malicious behaviors in smart contracts (Dwivedi et al., 2021; Boi et al., 2024; Hu et al., 2023; Wei et al., 2025; Chen et al., 2025; Sikder et al., 2025; Yu et al., 2025). It enables automated and intelligent auditing of smart contract code, reducing the cost of manual audits while improving audit efficiency and accuracy (Ajienka et al., 2020). In the future, as research deepens and technology advances, machine learning-based malicious code detection technology is certain to play an even greater role in the field of smart contracts.
@@ 2. Background
Smart contracts are self-executing and self-verifying computer protocols designed to facilitate, verify, or enforce the negotiation of contracts between parties (Barboni et al., 2023). Smart contracts enable trustworthy transactions without third-party intermediaries, thereby reducing the risks of fraud and default. The concept of smart contracts was first proposed by computer scientist Nick Szabo in 1994 (Ghaleb et al., 2023; Wu et al., 2025; Xiao et al., 2025; Cai et al., 2024; Han et al., 2026; Tas et al., 2024).
Smart contracts are one of the core applications of blockchain technology, implemented by deploying programmable scripts on the blockchain (Sharma et al., 2023; Aguilar et al., 2024; Wang et al., 2026). These scripts can automatically trigger, execute, and verify relevant operations when specific conditions are met (Tolmach et al., 2021). Since smart contracts run on decentralized blockchain networks, they have the following characteristics:
@@ 3.1. Feature Extraction
The features obtained from Ethereum smart contracts cannot be directly used as machine learning inputs, thus vectorization of smart contract features becomes a necessary step (Wood and others, 2014; Arceri et al., 2024). This paper employs an opcode-based text vectorization method that first categorizes opcodes according to their semantic meanings in the Ethereum Virtual Machine (EVM), then transforms the simplified opcodes into vectors using N-Gram and TF-IDF algorithms.
@@ 3.2. Problem Analysis
In recent years, machine learning-based smart contract vulnerability detection techniques have made significant progress. These methods analyze smart contract source code, bytecode, and opcodes to better detect potential vulnerabilities (Zhuang et al., 2020, 2021; Li et al., 2025a; Sun et al., 2025; Zenggang et al., 2025; He et al., 2025).
@@ 4.1. Dataset Collection
Etherscan (Tan Matai, ) is an Ethereum blockchain explorer and analytics platform. It allows users to query and retrieve information about transactions, addresses, blocks, tokens (including ERC-20 and ERC-721 tokens), and smart contracts on the Ethereum network. Etherscan provides developers, researchers, and other participants in the Ethereum ecosystem with a user-friendly interface to track and analyze activities on the Ethereum blockchain.
@@ 4.2. Label Annotation
We used the Ethereum static analysis tool Slither (Feist et al., 2019) to detect vulnerabilities in smart contracts. After configuring the Slither tool and selecting the compiler version (solc) according to the pragma solidity ˆ0.6.0 statement in the smart contract code, we performed the detection. The detection results for one sample smart contract (as shown in Figure 4) revealed both access control vulnerabilities and reentrancy attack vulnerabilities.
@@
(7) $\text{F1-score}=\frac{2\times\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$
@@ 6. Related Work
In machine learning-based malicious code detection, feature extraction and representation are crucial (Liu and Li, 2025). Researchers attempt to extract various features from smart contract code, including syntactic features, semantic features, control flow, and data flow (Bu et al., 2025b). These features can be represented using techniques such as the bag-of-words model, word vector embeddings, and graph representation learning (Asudani et al., 2023). Recent studies further demonstrate the importance of transaction pattern analysis, as seen in (Li et al., 2024c), which detects malicious Web3 accounts through transaction graph analysis. Effective feature extraction and representation methods help improve the detection performance of machine learning models (Li et al., 2024d).
@@
In 2021, Xu et al. (Xu et al., 2021) extracted corresponding abstract syntax trees from the source code of each vulnerability dataset, as well as from a large dataset. They then extracted feature vectors based on subnodes shared by the two ASTs. Finally, they used the K-Nearest Neighbor (KNN) algorithm and Stochastic Gradient Descent (SGD) to train models. Both methods leveraged abstract syntax trees as the basis for feature extraction, avoiding the need to set up complex patterns or execute them on Ethereum. Recent advancements like (Li et al., 2024e) and (Li et al., 2024b) further extend this paradigm by detecting state defects in decentralized exchanges and analyzing cross-contract execution patterns. This makes these methods relatively simple for professionals to use. However, a limitation of these approaches is that obtaining source code is more difficult compared to other contract attributes, reducing their practical utility for third-party detection. Alternative solutions like (Li et al., 2020) and (Liu et al., 2024b) address this challenge by characterizing account behaviors and detecting malicious activities through gas usage patterns.
@@
Research on machine learning-based malicious code detection in smart contracts has made progress but still faces many challenges and problems. The integration of AI and blockchain for privacy preservation (Li et al., 2023), the detection of emerging threats like NFT wash trading (Niu et al., 2024), and the need to characterize complex ecosystems like Solana NFTs (Kong et al., 2024) highlight the evolving nature of this field. Additionally, emerging tools like (Wang et al., 2024b) statistically quantify external data dependencies in real-world contracts, while (Li et al., 2025b) demonstrates how LLMs can automate bad practice detection, pointing to future research directions.